# BOTS: A Unified Framework for Bayesian Online Task Selection in LLM Reinforcement Finetuning

### v2

@@ Abstract
Reinforcement finetuning (RFT) is a key technique for aligning Large Language Models (LLMs) with human preferences and enhancing reasoning, yet its effectiveness is highly sensitive to which tasks are explored during training. Uniform task sampling is inefficient, wasting computation on tasks that are either trivial or unsolvable, while existing task selection methods often suffer from high rollout costs, poor adaptivity, or incomplete evidence. We introduce BOTS, a unified framework for Bayesian Online Task Selection in LLM reinforcement finetuning. Grounded in Bayesian inference, BOTS adaptively maintains posterior estimates of task difficulty as the model evolves. It jointly incorporates explicit evidence from direct evaluations of selected tasks and implicit evidence inferred from these evaluations for unselected tasks, with Thompson sampling ensuring a principled balance between exploration and exploitation. To make implicit evidence practical, we instantiate it with an ultra-light interpolation-based plug-in that estimates difficulties of unevaluated tasks without extra rollouts, adding negligible overhead. Empirically, across diverse domains and LLM scales, BOTS consistently improves data efficiency and performance over baselines and ablations, providing a practical and extensible solution for dynamic task selection in RFT.  [^1] Our code is released at https://github.com/modelscope/Trinity-RFT/tree/main/examples/bots
@@ 1 Introduction
Reinforcement finetuning (RFT) has become a key technique for aligning Large Language Models (LLMs) with human preferences and enhancing their reasoning capabilities (jaech2024openai; guo2025deepseek; luo2025deepscaler; hu2025open; zeng2025simplerl). However, the effectiveness of RFT is highly sensitive to task selection (parashar2025curriculum; shen2025skywork; zhu2025eduflow; wen2025structured; li2025temple). Naively training on a static, uniformly sampled dataset is inefficient: the model spends excessive computation on tasks that are either already mastered (too easy) or beyond reach (too hard) (yu2025dapo; bae2025online; chen2025self). This inefficiency not only inflates training costs but also destabilizes optimization by reducing the effective batch size. The central challenge, therefore, is to dynamically select tasks of “just right” difficulty to maximize learning efficiency as the model’s capability evolves.
Existing methods to this challenge face several limitations. Offline task selection (parashar2025curriculum; shen2025skywork; zhu2025eduflow; wen2025structured; li2025temple), which pre-schedules tasks from easy to hard, is too rigid and does not adapt to the evolving trajectory of the model. In response, a few online selection methods have been proposed, aiming to adaptively choose tasks based on model’s current capability. Core challenge of these methods lies in the tradeoff between the computational cost of collecting information and the accuracy of the resulting performance estimates. We argue that existing solutions are not sufficiently efficient: some expend excessive computation on information gathering, undermining efficiency, while others fail to fully exploit collected information, leading to suboptimal selection. On one hand, oversampling-based methods (yu2025dapo; bae2025online) find suitable tasks by rolling out oversized batches, introducing substantial extra cost. On the other hand, non-oversampling approaches typically rely on a single source of information—either leveraging historical evaluations as explicit evidence (chen2025self; qu2025can) or exploiting inter-task correlations as implicit evidence (sun2025improving). Our empirical results reveal a clear complementarity: explicit evidence provides stable and accurate task-difficulty estimates but suffers from a slow warm-up when historical evaluations are scarce in early training, whereas implicit evidence quickly guides early-stage selection yet becomes less reliable in later stages. These findings indicate that relying solely on one type of evidence leaves information underutilized and leads to suboptimal task selection. Therefore, a principled framework to fuse these complementary evidence sources is essential for robust and efficient online task selection.
@@
Figure 1: Overview of the BOTS framework. BOTS operates in a continuous loop of task selection, model training, and posterior updating. (1) Selection: Thompson sampling from the posterior beliefs selects a batch of tasks whose estimated success probabilities are near a target difficulty (e.g., $p^{*}=0.5$). (2) Training & Evidence Collection: The LLM is finetuned, yielding direct success/failure counts (explicit evidence) for the selected batch. For unselected tasks, predicted counts (implicit evidence) are produced by a plug-in; in Section 3.3, we introduce an ultra-lightweight interpolation-based variant with negligible overhead. (3) Posterior Updating: Explicit and implicit evidence are fused using our generalized Bayesian update rule (Section 3.2).
@@ 2 Related Works
The impact of task difficulty on RFT of LLMs has become an active research topic. Inspired by the seminal idea of curriculum learning (bengio2009curriculum), researchers have proposed various strategies for selecting appropriate tasks in LLM RFT. For instance, parashar2025curriculum; shen2025skywork advocate scheduling tasks from easy to hard, enabling LLMs to gradually acquire reasoning skills. Similar ideas have been extended to multi-modal LLMs (zhu2025eduflow; wen2025structured; li2025temple), though these methods mainly focus on offline task selection following an easy-to-hard trajectory.
More recently, online selection strategies have emerged, often targeting tasks of moderate difficulty. One line of work adopts sampling-based task filtering, where tasks with consistently trivial rewards (all zeros or ones) are considered uninformative and are down-weighted or filtered (yu2025dapo; bae2025online). While effective, these methods require additional rollouts, incurring non-trivial overhead. To avoid extra rollouts, several works attempt to predict task passing rates without direct rollouts. chen2025self formulate task selection as a non-stationary multi-armed bandit problem, treating each problem category (e.g., difficulty level or type) as an arm and using absolute advantage as a reward proxy, with posterior estimation based on historical outcomes. qu2025can extend this framework to the task level. Although these methods eliminate extra rollouts, they rely solely on direct evaluations and overlook cross-task relationships. In contrast, sun2025improving propose evaluating a small set of reference tasks and predicting the passing rates of others using an attention-inspired kernel over embeddings. However, this approach still requires additional rollouts for the reference set and discards historical evaluation information.
More comprehensive discussion and comparison on related works are provided in Appendix C.
@@ 3.2 Core Mechanism: Fusing Evidence in a Unified Posterior
where $\alpha_{0},\beta_{0}$ denote the prior parameter set for the Beta distribution, and the coefficient $\lambda\in[0,1]$ discounts historical information;
@@
The proof is placed in Appendix D.1.
@@ 3.3 Ultra-Light Interpolation Plug-in for Implicit Evidence
$\tilde{p}(\mathcal{B}_{t},k)=\mathrm{clip}\,\Big(\tilde{\mu}_{t}(\mathcal{B}_{t})\,\bar{p}_{s}^{k}+\bigl(1-\tilde{\mu}_{t}(\mathcal{B}_{t})\bigr)\,\bar{p}_{w}^{k},\;0,\;1\Big).$ (5)
Computational Cost. The primary cost of our approach arises from evaluating the reference models for all tasks, which requires extensive rollouts. Fortunately, benchmarking standard base models has already become routine in RL dataset construction—for example, cheng2025revisiting report Qwen2.5-7B-Instruct (Yang2024Qwen25TR) and Qwen3-30B-A3B (yang2025qwen3) performance as meta-tags for filtering. Given these pre-computed references, the additional online overhead of our interpolation-based estimator is negligible compared to training-time rollouts (see Appendix F.1 for empirical results). This efficiency follows from two properties: (i) no extra rollouts are needed during training, and (ii) both updates and predictions reduce to lightweight vector operations.
@@ 3.4 Thompson Sampling for Task Selection
The first question is: at what difficulty level does the current model benefit most from training? Prior works (chen2025self; sun2025improving) show that, under binary rewards, tasks with success probability around $0.5$ are most informative for learning as they lead to gradients with larger expected magnitude than tasks with success probability close to 0 or $1$. We define the utility of a task as the absolute deviation of its posterior mean from a target success probability $p^{*}\in(0,1)$, with $p^{*}=0.5$ as the canonical choice.
@@
Given this target success probability, the problem of online task selection naturally reduces to a non-stationary bandit problem. The central challenge is the tradeoff between exploitation and exploration: the model must decide whether to select tasks with high-confidence estimates close to the target rate in order to maximize immediate utility, or to select tasks with high uncertainty to gather information that may improve future decisions. A purely exploitative strategy might select tasks whose posterior mean $\tilde{p}_{t}^{k}$ is closest to $p^{*}$, but this risks overlooking tasks whose difficulty is currently uncertain but potentially optimal. To naturally balance exploration and exploitation, we employ Thompson Sampling (thompson1933likelihood), a strategy renowned for both its empirical effectiveness and theoretical guarantees. More specifically, at each selection step $t$, we perform the following: (1) Posterior Sampling: Draw a sample of the passing rate from its current posterior distribution $\hat{p}_{k}\sim\mathrm{Beta}(\alpha_{t}^{k},\beta_{t}^{k})$ for each task $\mathcal{T}^{k}$ in the pool. (2) Selection: Select tasks with the highest estimated utilities $\{\hat{u}_{k}:=|\hat{p}_{k}-p^{*}|\}$ to form the training batch $\mathcal{B}_{t+1}$. This procedure elegantly prioritizes tasks that are likely to be near the target difficulty, while the inherent variance in sampling from the posterior ensures that tasks with higher uncertainty are naturally explored.
@@
Impact of $\lambda,\rho$ on Task Selection. Proposition 3.4 characterizes how the hyperparameters $\lambda$ and $\rho$ jointly determine the effective sample size $n_{t}=\alpha_{t}+\beta_{t}$ of each task’s Beta posterior, thereby controlling the confidence of the estimated probability of success.
@@
The proof is given in Appendix D.2. Specifically, $\lambda$ controls the overall scale of the effective sample size: smaller values accelerate forgetting and increase posterior uncertainty, promoting exploration, while larger values stabilize the posterior and favor exploitation. Meanwhile, $\rho$ sets the lower bound of effective counts for unevaluated tasks: smaller values maintain high uncertainty to encourage exploration, whereas larger values reduce uncertainty and bias selection toward exploitation. Together, $\lambda$ and $\rho$ provide complementary knobs for balancing exploration and exploitation in online task selection. A more detailed discussion is placed in Appendix E.1.
@@ 4 Experiments
We begin with Section 4.1, which introduces datasets, reinforcement finetuning protocols, evaluation metrics, and computational cost. Section 4.2 and Section 4.3 analyze the effects of mixing evidence ($\rho$) and forgetting/adaptivity ($\lambda$) on task selection respectively. Section 4.4 then compares BOTS with competitive baselines across model scales and domains, while Section 4.5 summarizes additional experiments provided in the Appendix F.
@@ 4.1 Setups
Dataset.  We conduct experiments on GURU (cheng2025revisiting), a well-curated cross-domain RL dataset. Each subset is deduplicated, verified, and filtered. We use its math, code, and logic subsets (excluding the Zebra Puzzle due to its non-binary reward). Detailed information about the used datasets is provided in Appendix B.1 (for training) and Appendix B.2 (for evaluation).
RFT Setting.  We adopt GRPO (grpo2024), Qwen2.5-1.5B-Instruct and Qwen2.5-7B. Key hyperparameters include a learning rate of 1e-6, 16 rollouts per task, and a temperature of $1.0$. Comprehensive training details and used RL algorithm are provided in Appendix B.3 and B.5.
Evaluation.  We report the following metrics to evaluate our framework and its ablations, with formal definitions given in Appendix B.4.
@@
• Best-so-far (BSF).  It measures the performance gain relative to the random baseline under a fixed budget. Within specific ratio (25%, 50%, 100%) of total training steps, BSF is the ratio between a method’s best-so-far performance and the baseline’s best-so-far performance. For example, at step 50 (total steps 100), if the baseline’s best is 0.4 and a method’s best is 0.6, then BSF(50%) = $0.6/0.4=1.5$. The baseline always has BSF = 1; larger values indicate greater gains.
Computational Overhead. We note that BOTS introduces negligible additional computation, with overhead measured at $\leq 0.2\%$ of total training time (see analysis in Section 3.3 and empirical results in Appendix F.1). Thus, in the main results, we report TTB and BSF, in terms of training steps rather than wall-clock time, as their difference is practically insignificant.
4.2 Fusing Two sources of Evidence: Analyzing the Impact of $\rho$
Recall as defined in Equation 1, $\rho$ controls the relative weights of explicit and implicit evidence. A smaller $\rho$ places greater emphasis on explicit evidence, whereas a larger $\rho$ increases the contribution of implicit evidence. Moreover, according to Proposition 3.4 and the analysis in Section 3.4, $\rho$ determines the limit inferior of the equivalent counts in the Beta posterior. Intuitively, reducing $\rho$ increases the overall randomness of posterior sampling, while enlarging $\rho$ stabilizes the estimates by anchoring them more strongly to implicit evidence.
We investigate the role of $\rho$ in task selection by varying $\rho\in\{0.0,0.05,0.1,0.2,0.5,1.0\}$, while keeping other hyperparameters fixed at their default values ($\lambda=0.1$, posterior sampling enabled).
Benchmark MATH500 AIME24 Math Aggregated Performance
$\lambda=0.1,\rho=0.0$ 1.10 0.73 0.88 1.00 0.96 1.03 0.57 0.92 - 1.00 1.75 0.92 1.13 0.73 0.99 1.02 0.98 1.01
$\mathbf{\lambda=0.1,\rho=0.05}$ 1.02 0.56 0.60 1.08 1.06 1.05 0.50 0.51 0.68 1.00 2.50 1.17 1.07 0.58 0.60 1.10 1.05 1.03
$\mathbf{\lambda=0.1,\rho=0.1}$ 0.89 0.49 0.57 1.13 1.05 1.05 0.51 1.00 1.00 1.25 1.75 1.00 0.89 0.56 0.64 1.12 1.07 1.05
$\mathbf{\lambda=0.1,\rho=0.2}$ 0.99 0.68 0.68 1.10 1.10 1.06 0.96 1.06 - 1.25 1.50 0.92 0.99 0.74 0.80 1.13 1.08 1.08
$\lambda=0.1,\rho=0.5$ 0.93 0.64 0.67 1.10 1.10 1.03 0.63 1.11 1.12 0.75 2.25 1.00 0.93 0.66 1.00 1.10 1.11 1.01
$\lambda=0.1,\rho=1.0$ 0.92 0.61 0.95 1.04 1.08 1.02 0.42 0.71 - 1.25 2.25 0.92 0.90 0.61 0.86 1.05 1.06 1.01
Table 1: TTB and BSF evaluated on MATH500, AIME24, and aggregated performance with Qwen2.5-1.5B-Instruct. For TTB, notation “-” indicates that the target performance is never achieved within the evaluation window. The best and second best results are marked accordingly.

### v3

@@ Abstract
Reinforcement finetuning (RFT) is a key technique for aligning Large Language Models (LLMs) with human preferences and enhancing reasoning, yet its effectiveness is highly sensitive to which tasks are explored during training. Uniform task sampling is inefficient, wasting computation on tasks that are either trivial or unsolvable, while existing task selection methods often suffer from high rollout costs, poor adaptivity, or incomplete evidence. We introduce BOTS, a unified framework for Bayesian Online Task Selection in LLM reinforcement finetuning. Grounded in Bayesian inference, BOTS adaptively maintains posterior estimates of task difficulty as the model evolves. It jointly incorporates explicit evidence from direct evaluations of selected tasks and implicit evidence inferred from these evaluations for unselected tasks, with Thompson sampling ensuring a principled balance between exploration and exploitation for task selection. To make implicit evidence practical, we instantiate it with an ultra-light interpolation-based plug-in that estimates difficulties of tasks without extra rollouts, adding negligible overhead. Empirically, across diverse domains and LLM scales, BOTS consistently improves data efficiency and performance over baselines and ablations, providing a practical and extensible solution for dynamic task selection in RFT [^1] Our code is released at https://github.com/modelscope/Trinity-RFT/tree/main/examples/bots.
@@ 1 Introduction
Reinforcement finetuning (RFT) has become a key technique for aligning Large Language Models (LLMs) with human preferences and enhancing their reasoning capabilities (Jaech et al., 2024; Guo et al., 2025; Luo et al., 2025; Hu et al., 2025; Zeng et al., 2025). However, the effectiveness of RFT is highly sensitive to task selection (Parashar et al., 2025; Shen et al., 2025; Zhu et al., 2025; Wen et al., 2025; Li et al., 2025a). Naively training on a static, uniformly sampled dataset is inefficient: the model spends excessive computation on tasks that are either already mastered (too easy) or beyond reach (too hard) (Yu et al., 2025; Bae et al., 2025; Chen et al., 2025b). This inefficiency not only inflates training costs but also destabilizes optimization by reducing the effective batch size. The central challenge, therefore, is to dynamically select tasks of “just right” difficulty to maximize learning efficiency as the model’s capability evolves.
Existing methods to this challenge face several limitations. Offline task selection (Parashar et al., 2025; Shen et al., 2025; Zhu et al., 2025; Wen et al., 2025; Li et al., 2025a), which pre-schedules tasks from easy to hard, is too rigid and does not adapt to the evolving trajectory of the model. In response, a few online selection methods have been proposed, aiming to adaptively choose tasks based on model’s current capability. Core challenge of these methods lies in the tradeoff between the computational cost of collecting information and the accuracy of the resulting performance estimates. We argue that existing solutions are not sufficiently efficient: some expend excessive computation on information gathering, undermining efficiency, while others fail to fully exploit collected information, leading to suboptimal selection. On one hand, oversampling-based methods (Yu et al., 2025; Bae et al., 2025) find suitable tasks by rolling out oversized batches, introducing substantial extra cost. On the other hand, non-oversampling approaches typically rely on a single source of information—either leveraging historical evaluations as explicit evidence (Chen et al., 2025b; Qu et al., 2025) or exploiting inter-task correlations as implicit evidence (Sun et al., 2025). Our empirical results reveal a clear complementarity: explicit evidence provides stable and accurate task-difficulty estimates but suffers from a slow warm-up when historical evaluations are scarce in early training, whereas implicit evidence quickly guides early-stage selection yet becomes less reliable in later stages. These findings indicate that relying solely on one type of evidence leaves information underutilized and leads to suboptimal task selection. Therefore, a principled framework to fuse these complementary evidence sources is essential for robust and efficient online task selection.
@@
[…]
@@ 2 Related Works
The impact of task difficulty on RFT of LLMs has become an active research topic. Inspired by the seminal idea of curriculum learning (Bengio et al., 2009), researchers have proposed various strategies for selecting appropriate tasks in LLM RFT. For instance, Parashar et al. (2025); Shen et al. (2025) advocate scheduling tasks from easy to hard, enabling LLMs to gradually acquire reasoning skills. Similar ideas have been extended to multi-modal LLMs (Zhu et al., 2025; Wen et al., 2025; Li et al., 2025a), though these methods mainly focus on offline task selection following an easy-to-hard trajectory.
More recently, online selection strategies have emerged, often targeting tasks of moderate difficulty. One line of work adopts sampling-based task filtering, where tasks with consistently trivial rewards (all zeros or ones) are considered uninformative and are down-weighted or filtered (Yu et al., 2025; Bae et al., 2025). While effective, these methods require additional rollouts, incurring non-trivial overhead. To avoid extra rollouts, several works attempt to predict task passing rates without direct rollouts. Chen et al. (2025b) formulate task selection as a non-stationary multi-armed bandit problem, treating each problem category (e.g., difficulty level or type) as an arm and using absolute advantage as a reward proxy, with posterior estimation based on historical outcomes. Qu et al. (2025) extend this framework to the task level. Although these methods eliminate extra rollouts, they rely solely on direct evaluations and overlook cross-task relationships. In contrast, Sun et al. (2025) propose evaluating a small set of reference tasks and predicting the passing rates of others using an attention-inspired kernel over embeddings. However, this approach still requires additional rollouts for the reference set and discards historical evaluation information.
More comprehensive discussion and comparison on related works are provided in Appendix E.
@@ 3.2 Core Mechanism: Fusing Evidence in a Unified Posterior
where $\alpha_{0},\beta_{0}$ denote the prior parameter set for the Beta distribution, and the coefficient $\lambda\in[0,1]$ discounts historical information by interpolating the counts with the prior (Raj and Kalyani, 2017);
@@
The proof is placed in Appendix B.1.
@@ 3.3 Ultra-Light Interpolation Plug-in for Implicit Evidence
$\tilde{p}(k,\mathcal{B}_{t})=\mathrm{clip}\,\Big(\tilde{\mu}_{t}(\mathcal{B}_{t})\,\bar{p}_{s}^{k}+\bigl(1-\tilde{\mu}_{t}(\mathcal{B}_{t})\bigr)\,\bar{p}_{w}^{k},\;0,\;1\Big).$ (5)
@@ 3.4 Thompson Sampling for Task Selection
The first question is: at what difficulty level does the current model benefit most from training? Prior works (Chen et al., 2025b; Sun et al., 2025) show that, under binary rewards, tasks with success probability around $0.5$ are most informative for learning as they lead to gradients with larger expected magnitude than tasks with success probability close to 0 or $1$. We define the utility of a task as the absolute deviation of its posterior mean from a target success probability $p^{*}\in(0,1)$, with $p^{*}=0.5$ as the canonical choice.
@@
Given this target success probability, the problem of online task selection naturally reduces to a non-stationary bandit problem. The central challenge is the tradeoff between exploitation and exploration for task selection: the model must decide whether to select tasks with high-confidence estimates close to the target rate in order to maximize immediate utility, or to select tasks with high uncertainty to gather information that may improve future decisions. A purely exploitative strategy might select tasks whose posterior mean $\tilde{p}_{t}^{k}$ is closest to $p^{*}$, but this risks overlooking tasks whose difficulty is currently uncertain but potentially optimal. To naturally balance exploration and exploitation, we employ Thompson Sampling (Thompson, 1933), a strategy renowned for both its empirical effectiveness and theoretical guarantees. More specifically, at each selection step $t$, we perform the following: (1) Posterior Sampling: Draw a sample of the passing rate from its current posterior distribution $\hat{p}_{k}\sim\mathrm{Beta}(\alpha_{t}^{k},\beta_{t}^{k})$ for each task $\mathcal{T}^{k}$ in the pool. (2) Selection: Select tasks with the highest estimated utilities $\{\hat{u}_{k}:=-|\hat{p}_{k}-p^{*}|\}_{k=1}^{N}$ to form the training batch $\mathcal{B}_{t+1}$. This procedure elegantly prioritizes tasks likely to be near the target difficulty, while the inherent variance in sampling from the posterior ensures that tasks with higher uncertainty are naturally explored.
3.5 Hyperparameters
The hyperparameters $\lambda$ and $\rho$ play distinct but complementary roles in shaping both the difficulty estimation process and the uncertainty profile that drives Thompson Sampling–based task selection. Their effects can be summarized along two axes:
Impact on difficulty estimation.
The parameter $\lambda$ controls the balance between adaptivity and stability in Bayesian belief updating. A larger $\lambda$ places more weight on recent evaluations, enabling the posterior to rapidly adapt to the model’s evolving capability. However, this increased adaptivity also makes the estimate more sensitive to noise in the most recent observations, reducing stability. Conversely, a smaller $\lambda$ assigns greater weight to accumulated history, reducing variance in the estimated probability of success. This leads to more stable but less responsive difficulty estimates. A toy example illustrating these effects is provided in Figure 4.
The parameter $\rho$ governs how explicit and implicit evidence are fused. A larger $\rho$ biases the update toward implicit evidence, causing tasks without direct evaluations to inherit stronger difficulty signals inferred from tasks with direct evaluations. A smaller $\rho$, in contrast, favors explicit evidence, making the posterior rely primarily on direct observations and thus behave more conservatively when a task has not been recently evaluated.
Impact on selection-time uncertainty.
@@ 3.5 Hyperparameters
Proposition 3.5 characterizes how the hyperparameters $\lambda$ and $\rho$ jointly determine the effective sample size $n_{t}=\alpha_{t}+\beta_{t}$ of each task’s Beta posterior, thereby controlling the confidence of the estimated probability of success.
@@
The proof is given in Appendix B.2. In terms of balancing exploration toward tasks whose current success probability estimates deviate from the target and exploitation of tasks estimated to be near the target success probability, $\lambda$ controls the overall scale of the effective sample size $n_{t}=\alpha_{t}+\beta_{t}$, which directly dictates the variance of the posterior distribution. A larger $\lambda$ accelerates discounting of older evidence, reducing $n_{t}$ and increasing posterior uncertainty. Under Thompson Sampling, this results in more exploratory behavior, particularly toward tasks whose current success-probability estimates deviate from the target difficulty. A smaller $\lambda$ enlarges $n_{t}$, decreasing posterior variance and nudging the strategy toward exploitation of tasks estimated to be near the target success probability.
The parameter $\rho$ primarily affects the effective sample sizes of tasks that have not been directly evaluated at the current step. A larger $\rho$ increases these tasks’ effective sample sizes via stronger pseudo-count updates, reducing their posterior uncertainty and making them less likely to be explored. A smaller $\rho$ decreases their effective sample sizes, increasing posterior variance and allowing tasks with difficulty estimates farther from the target to be selected, thereby improving long-run exploration.
3.6 Practical Considerations
Hyperparameters.
We recommend using $\lambda=0.1$ and $\rho=0.1$ as the default configuration for BOTS. This setting consistently performs well across all model scales and task domains evaluated in Section 4. For new applications, practitioners may start from this default and subsequently adjust the hyperparameters using the insights provided in Section 3.5.
Computational Cost of Interpolation-based Task Difficulty Estimation.
The main cost of the interpolation-based estimator (Section 3.3) comes from evaluating reference models on all tasks. Fortunately, such evaluations are increasingly common in modern RL datasets—for example, the GURU dataset (Cheng et al., 2025) provides Qwen2.5-7B-Instruct (Yang et al., 2024) and Qwen3-30B-A3B (Yang et al., 2025) scores as difficulty tags for offline filtering. Our estimator simply reuses these tags for online task selection. In settings where reference-model evaluations are not available, enabling implicit evidence requires a one-time rollout cost to compute them. Given the substantial gains that implicit evidence brings, this introduces a natural computation–efficiency tradeoff. Unlike oversampling-based online methods, this cost does not recur during training and amortizes well when training multiple models or checkpoints. Additional practical benefits of this computation are discussed in Appendix F.3. Once the reference scores are available, the online overhead of the interpolation-based estimator is negligible (see Appendix G.1), since no extra rollouts are required and all updates reduce to simple vector operations.
Choice of Reference Models.
When a dataset provides pre-computed base model evaluation tags, as in GURU, we recommend directly using these tags as reference-model signals. This avoids any additional rollout cost, even though it may introduce extrapolation (e.g., when both reference models are stronger or weaker than the training model). Our empirical results in Appendix G.2 show that while extrapolation can reduce the accuracy of implicit evidence, BOTS remains robust, making the computational savings generally worthwhile. When multiple reference-model pairs are available, we suggest choosing models with a clear capability gap, which provides stronger discrimination for interpolating task difficulty. For instance, GURU’s choice of Qwen2.5-7B-Instruct and Qwen3-30B-A3B yields informative difficulty signals across a wide range. Finally, although the interpolation-based estimator supports extrapolation in principle, our results show that avoiding extrapolation produces more accurate implicit evidence and slightly stronger overall BOTS performance.
@@ 4 Experiments
We begin with Section 4.1, which introduces datasets, reinforcement finetuning protocols, evaluation metrics, and computational cost. Section 4.2 and Section 4.3 analyze the impact of hyperparameters ($\rho$), ($\lambda$) on task selection respectively. Section 4.4 then compares BOTS with competitive baselines across model scales and domains, while Section 4.5 summarizes additional experiments provided in the Appendix G.
@@ 4.1 Setups
Dataset.  We conduct experiments on GURU (Cheng et al., 2025), a well-curated cross-domain RL dataset. Each subset is deduplicated, verified, and filtered. We use its math, code, and logic subsets (excluding the Zebra Puzzle due to its non-binary reward). Detailed information about the used datasets is provided in Appendix D.1 (for training) and Appendix D.2 (for evaluation).
RFT Setting.  We adopt GRPO (Shao et al., 2024), Qwen2.5-1.5B-Instruct and Qwen2.5-7B. Key hyperparameters include a learning rate of 1e-6, 16 rollouts per task, and a temperature of $1.0$. Comprehensive training details and used RL algorithm are provided in Appendix D.3 and D.5.
Evaluation.  We report the following metrics to evaluate our framework and its ablations, with formal definitions given in Appendix D.4.
@@
• Best-so-far (BSF).  It measures the performance gain relative to the random baseline under a fixed budget. Within a specific ratio (25%, 50%, 100%) of total training steps, BSF is the ratio between a method’s best-so-far performance and the baseline’s best-so-far performance. For example, at step 50 (total steps 100), if the baseline’s best is 0.4 and a method’s best is 0.6, then BSF(50%) = $0.6/0.4=1.5$. The baseline always has BSF = 1; larger values indicate greater gains.
Computational Overhead. We note that BOTS introduces negligible train-time additional computation, with overhead measured at $\leq 0.2\%$ of total training time (see analysis in Section 3.3 and empirical results in Appendix G.1). Thus, in the main results, we report TTB and BSF, in terms of training steps rather than wall-clock time, as their difference is practically insignificant.
4.2 Analyzing the Impact of $\rho$
In Section 3.5, we analyzed the role of the hyperparameter $\lambda$ and its influence on BOTS. In this section, we empirically validate these analyses. We investigate the role of $\rho$ in task selection by varying $\rho\in\{0.0,0.05,0.1,0.2,0.5,1.0\}$, while keeping other hyperparameters fixed at their default values ($\lambda=0.1$, posterior sampling enabled).
Figure 2: Qwen2.5-1.5B-Instruct on Math with $\lambda=0.1,\rho\in\{0.0,0.05,0.1,0.2,0.5,1.0\}$. Ratio of sampled training tasks (measured over 16 rollouts) with passing rates: strictly between 0 and 1 (L1), strictly greater than 0 (L2), and strictly less than 1 (L3), along with aggregated performance (MATH500, AMC23 and AIME24), plotted against training steps. All results are averaged over 3 random seeds, and 95% confidence intervals are reported. More detailed results are placed in Appendix G.6.
Benchmark MATH500 AMC23 AIME24 Math Aggregated Performance
Target Fraction 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
$\rho=0.0$ 0.95 0.88 1.04 1.00 1.01 1.01 1.13 0.98 - 0.98 0.99 0.98 1.16 0.78 0.76 0.93 1.18 1.13 0.98 0.94 - 0.99 1.02 0.98
$\rho=0.05$ 0.77 0.66 0.66 1.05 1.08 1.04 0.86 0.78 0.72 1.11 1.08 1.12 0.66 0.77 0.61 1.20 1.05 1.32 0.78 0.68 0.64 1.06 1.09 1.06
$\rho=0.1$ 0.86 0.66 0.78 1.05 1.09 1.06 0.86 0.68 0.74 1.10 1.16 1.06 0.86 0.85 0.50 1.20 1.25 1.23 0.85 0.66 0.72 1.06 1.12 1.05
$\rho=0.2$ 0.76 0.61 0.74 1.08 1.08 1.04 0.87 0.62 0.70 1.05 1.20 1.04 1.00 0.71 0.47 1.02 1.63 1.31 0.78 0.63 0.69 1.07 1.10 1.05
$\rho=0.5$ 0.82 0.68 0.98 1.07 1.04 1.01 0.78 0.61 - 1.14 1.16 1.00 0.70 0.80 0.67 1.16 1.24 1.16 0.79 0.64 0.89 1.09 1.09 1.00
$\rho=1.0$ 0.78 0.70 - 1.06 1.04 0.98 0.85 0.86 - 1.04 1.13 1.00 0.96 0.66 0.48 1.02 1.43 1.22 0.78 0.69 0.99 1.05 1.07 1.01
Table 1: Qwen2.5-1.5B-Instruct on Math with $\lambda=0.1,\rho\in\{0.0,0.05,0.1,0.2,0.5,1.0\}$. For TTB, notation “-” indicates that the target performance is never achieved within the evaluation window. The best and second best results are marked accordingly.